A man is 4 times as old as his son. Five years ago, the sum of their ages was 30.Form an equation in a variable of your choice and hence calculate their present ages.
Answer by Maths68(1474)

A man is 4 times as old as his son. Five years ago, the sum of their ages was 30.Form an equation in a variable of your choice and hence calculate their present ages.
Let
Present age of Man = m
Present age of Son = s


Man is 4 times as old as his son.
m=4s..........(1)
Five years ago,
m-5=s-5
the sum of their ages was 30
(m-5)+(s-5)=30
m-5+s-5=30
m+s-10=30
m+s=30+10
m+s=40.............(2)
Put the value of m from (1) to (2)
4s+s=40
5s=40
5s/5=40/5
s=8
Put the value of s in (1)
m=4s
m=4(8)
m=32
Present age of Man = m = 32 years old
Present age of Son = s = 8 years old